OpenStudy (anonymous):

need help determine whether the graphs of the equations are parallel lines x=2 x=4

OpenStudy (anonymous):

yes they are parallel if you graph them.

OpenStudy (anonymous):

the slope of the line x=2 is? the slope of the line x=4 is

OpenStudy (anonymous):

their parallel to each other yes.

OpenStudy (anonymous):

ok but how do i determine what the slopes are for each

OpenStudy (anonymous):

gradient = rise/run therefore, the gradient for x=2 is 0/2=0 and the gradient for x=4 is 0/4 = 0 hence both lines are parallel

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x%3D2+and+x%3D4 as you can see in the graph the lines are parallel to each other

OpenStudy (anonymous):

so then the slopes would be?

OpenStudy (anonymous):

0